1. Consider the utility maximization problem max ?^? + ? s.t. ?? + ? = ? (i.e., the price of good ? is 1), where the constants ?, ?, and ? are positive, and the constant ? ∈ (0,1).

A. Find the (Marshallian) demand functions, ?∗(?,?) and ?∗(?,?).

B. Find the partial derivatives of the demand functions w.r.t. ? and ?, and check their signs.

C. Set ? = 1/2. What are the demand functions in this case? What is ?∗? Denote the maximal utility as a function of ? and ? as ?∗(?,?), the value function, also called the indirect utility function. Verify that ??∗ ?? ⁄ = −?∗(?,?).   
Answer #1

A. In order to find Marshallian Demand Functions, we maximize the utility function subject to the given budget constraint.

B. Only the partial derivative of y w.r.t p is negative. Others are positive. Change in x as a change in m is zero.

C. Indirect utility function shows the maximal utility.a-l 1. May u=x + y s.t. Paty=m (A) The Lagrangran can be Escitten as, L = x y tam by : al = axa-1_XP=0 ax =P al = 1-7=0 A=1.a aa- -C) a = 1½ , the demand fills a become - x=1P 1 -1 - 1ap) - 2 y m-) = m+ (ap3) n- Tap2 = m - 4 > yah m-#1 21. = 422) +verify that au*- x*(pm). Hence, verified.